OpenStudy (anonymous):

What is the orbital radius and speed of a synchronous satellite which orbits the earth once every 24h? Take G=6.67×10−11Nm2/kg2 , Mass of the earth is 5.98×1024 kg

OpenStudy (praxer):

$$t^2=\frac{2 \pi r^3}{GM_e}$$ $$r= \sqrt{\frac{t^2GM_e}{2\ pi}}$$ once you find r find v by the fomual $$v= \sqrt{\frac{GM_e}{r}}$$

OpenStudy (praxer):

reference to the first formula. for the second formula it is cube root not square root... #apologies.

OpenStudy (praxer):

@micheallusly did you get it..:)

OpenStudy (wolf1728):

Here are the formulas: